OpenStudy (anonymous):

Two systems of equations are shown below: System A 2x + y = 5 -4x + 6y = -2 System B -10x + 19y = -1 -4x + 6y = -2 Which of the following statements is correct about the two systems of equations? A.) They will have the same solutions because the first equation of System B is obtained by adding the first equation of System A to 2 times the second equation of System A. B.) They will have the same solution because the first equation of System B is obtained by adding the first equation of System A to 3 times the second equation of System A. C.) The value of x for System B will be –5 times the value of x for System A because the coefficient of x in the first equation of System B is –5 times the coefficient of x in the first equation of System A. D.) The value of x for System A will be equal to the value of y for System B because the first equation of System B is obtained by adding –12 to the first equation of System A and the second equations are identical.

OpenStudy (cwrw238):

well carry out the instructions in each of the options to see which one fits eg try option 1: 2 times 2nd equation in system A = -8x + 12y = -4 add the first equation in system A -8x + 12y = -4 + 2x + y = 5 = -6x + 13y = 1 which is not the first equation in System B so option 1 is not correct